# Monodromy in the space of symmetric cubic surfaces with a line

### v1

@@ 5. Calculating the monodromy group
In this section, we will determine the monodromy group of the cover $\widetilde{\mathcal{S}}\to\mathcal{S}$ by a combination of classical, moduli-theoretic, and computational techniques. We begin with the following basic fact:
@@
5.1. Stability of the $D_{8}$ tritangent
We can further restrict the symmetric monodromy via an understanding of how $S_{4}$ acts on the 27 lines on a symmetric cubic surface. The following result of the first named author proves that it is independent of the choice of smooth symmetric cubic surface.
@@ 5.1. Stability of the $D_{8}$ tritangent
Once labels are fixed on the 27 lines, we can construct $W(E_{6})$ as a permutation group, given as the adjacency-preserving permutations of the 27 lines. As a subgroup of $S_{27}$ with the labeling of the lines coming from the Fermat cubic surface, the generators for $W$ are listed in data A.2.
@@ 5.2. Cycle decompositions in $W(E_{6})$
Proposition 5.11.
Let $g\in N_{W(E_{6})}(S_{4})\smallsetminus S_{4}$, let $\ell$ be a line on a symmetric cubic surface $X$, and let $\sigma\in S_{4}$. Then
$g(\sigma\ell)=\sigma(g\ell).$
Proof.
By 2.9, any such element $g$ centralizes $S_{4}$ within $W(E_{6})$, so $\sigma g$ and $g\sigma$ act identically on any line $\ell\subset X$. ∎
@@
Let $g$ be any element in the monodromy group, and let $\ell$ be a line on a symmetric cubic surface $X$. Then
  1. (1)
    If $g$ fixes $\ell$, then $g$ fixes $\sigma\ell$ for all $\sigma\in S_{4}$.
  2. (2)
    If $g\ell\neq\ell$ then $g(\sigma\ell)\neq g(\sigma\ell)$ for all $\sigma\in S_{4}$.
In particular the monodromy group acts on entire orbits simultaneously.
Proof.
If $g\in S_{4}$ this is clear. If $g\notin S_{4}$ this follows from 5.11. ∎
Proposition 5.13.
The symmetric monodromy action does not change the isotropy group of any line. Phrased differently, it acts on each $S_{4}$-orbit of lines independently.
Proof.
For $D_{8}$ this is clear since symmetric monodromy stabilizes each line, so they remain in a $D_{8}$ orbit. To see that the $S_{4}/C_{2}^{o}$ and $S_{4}/C_{2}^{e}$ orbits cannot be interchanged by the monodromy action, it suffices to observe that $C_{2}^{o}$ and $C_{2}^{e}$ are not conjugate in $W(E_{6})$. ∎
Lemma 5.14.
If the monodromy group contains two distinct non-trivial elements (i.e. if it is not trivial or cyclic of order two), then it is the Klein 4-group
$K_{4}\cong\left\langle\sigma_{1}\tau_{2},\tau_{1}\right\rangle.$
Proof.
We can look through the 16 elements of $\left\langle\sigma_{1},\sigma_{2},\tau_{1},\tau_{2}\right\rangle$ and ask whether each element satisfies the necessary constraints to lie in the symmetric monodromy group.
Since each element in the larger group is an involution, it is a product of disjoint transpositions, the number of which is well-defined in this monodromy problem (since any other choice of basepoint for monodromy conjugates the permutation representation of the monodromy class by an element in $W(E_{6})$, but the transposition length of an involution is clearly independent under such a conjugation). We can eliminate the following elements directly for having transposition length 10:
$\sigma_{2}\tau_{1},\tau_{2},\ \sigma_{1}\sigma_{2},\ \sigma_{1},\ \sigma_{1}\sigma_{2}\tau_{1}\tau_{2},\ \tau_{1}\tau_{2},\ \sigma_{2}.$
The remaining non-identity elements have transposition length 6 or 12, so we can ask whether they fix or permute each element of the two orbits. We can eliminate the following three elements, all of transposition length six, since they violate 5.12:
$\tau_{2},\ \sigma_{2}\tau_{2},\ \sigma_{1}\sigma_{2}\tau_{2}.$
Therefore our desired monodromy group $M$ is contained in what is left over, meaning we have a subset inclusion
$M\subseteq\left\{\operatorname{id},\ \tau_{1},\ \sigma_{2}\tau_{1},\ \sigma_{1}\tau_{2},\ \sigma_{1}\tau_{1},\ \sigma_{1}\tau_{1}\tau_{2},\ \sigma_{1}\sigma_{2}\tau_{1}\right\}.$
If the monodromy group is non-trivial, then it might be cyclic of order two generated by any one non-identity element in this set. The other case is that it contains two non-identity elements, in which case it must also include their product. In this case, we can exclude those elements in the set whose product with any non-identity element lands outside the set. These elements are $\sigma_{2}\tau_{1}$, $\sigma_{1}\tau_{1}$, and $\sigma_{1}\sigma_{2}\tau_{1}$. We are left with the Klein 4-group
$K_{4}=\left\{\operatorname{id},\ \tau_{1},\ \sigma_{1}\tau_{2},\ \sigma_{1}\tau_{1}\tau_{2}\right\}.$
These elements are explicitly given as permutations in data A.5. ∎
5.3. Certified tracking
Via 5.14, our computation of the monodromy group reduces to demonstrating the existence of two loops in the moduli space of smooth symmetric cubic surfaces which induce different permutations on the 27 lines. In conversations with T. Brysiewicz, working with his Pandora software [Bry24], we were able to generate strong evidence that this is true, and determine candidate loops.
While homotopy continuation software can generate strong evidence towards a computation, more refined algorithms are needed to turn these computations into proof. At each stage of tracking solutions along a one-parameter path, a guarantee is needed that paths don’t collide, and therefore that the computed permutation is indeed correct. These more sophisticated (and time-costly) methods are called certified tracking algorithms. Recent work of T. Duff and K. Lee provides algorithms which, among other things, are applicable for certifying computations in monodromy, bridging the gap between computation and proof [DL24, Theorem 1]. In conversations with Lee, their software is able to mathematically certify the following result.
Lemma 5.15 (Numerical certification).
There are two loops in the symmetric locus which induce distinct non-identity permutations on the set of $27$ lines.
We can now prove Theorem 1.2:
Proof of Theorem 1.2.
5.14 tells us that the monodromy group $M$ is contained in a specific $K_{4}<W(E_{6})$. By 5.15, there are two loops in the symmetric locus which generate distinct non-trivial elements in the monodromy group. Thus these two loops are a generating set for $K_{4}$. ∎
5.4. The incidence variety of 27 lines over the symmetric locus
Now that we have determined the monodromy group of the cover $\widetilde{\mathcal{S}}\to\mathcal{S}$ is $K_{4}$, we can actually say more about the topology of the space of symmetric cubic surfaces with a line:
Corollary 5.16.
The incidence variety of 27 lines restricted to the symmetric locus has 12 connected components. Explicitly as a $K_{4}$-set, the fiber over any symmetric cubic surface is of the form
@@ 5.4. The incidence variety of 27 lines over the symmetric locus
To conclude, we have an equivariant analog of 2.8:
@@
The space $\widetilde{\mathcal{S}}$ is a naturally a (disconnected) complex manifold. Moreover, the equivariant monodromy representation
$\pi_{1}(\mathcal{S},X)\to\operatorname{Aut}_{S_{4}}(H^{2}(X,\mathbb{Z}),\eta_{X})\cong S_{4}\times K_{4}$
is not surjective and has image isomorphic the Klein $4$-group $K_{4}$.
@@ 6.3. Our groups are not reflection groups
Suppose for the sake of contradiction that $K_{4}$ was a reflection subgroup; it would then take on the form of $W(A_{1})\times W(A_{1})$. Since each of the generators $s_{i}$ act on the 27 lines as a product of six disjoint transpositions, there are two nontrivial elements of $W(A_{1})\times W(A_{1})$ that are the product of six disjoint transpositions. However, data A.5 tells us that the symmetric monodromy group only has one element that is the product of six disjoint transpositions, a contradiction. ∎
@@ 6.4. Symmetric monodromy in the projective orthogonal groups
Recall that the set of six skew lines $\{\ell_{1},\ell_{3},\ell_{10},\ell_{11},\ell_{16},\ell_{22}\}$ determine a marking of the homology of $S$, where each line corresponds to the homology classes of orthogonal $(-1)$-exceptional curves $e_{1},e_{2},e_{3},e_{4},e_{5},e_{6}$ on $S$. Using data A.5, we can calculate how $K_{4}=\langle\tau_{1},\sigma_{1}\tau_{2}\rangle$ acts on the exceptional $(-1)$-curves, which in turns explicitly determines how $K_{4}$ acts on the $E_{6}$ lattice. Then by passing to the root lattice quotient used in the proof of the exceptional isomorphism outlined in 3.3, this $K_{4}$ projects to the symmetric monodromy group $K_{4}$ inside of $\operatorname{PO}(4,1,\mathbb{F}_{3})$.
@@
As Beauville remarks [Bea09, pg. 19], what makes this difficult is that it is unknown how to produce a marking of a cubic surface from a framing of the corresponding cyclic cubic 3-fold. A resolution to this problem would shed further light on symmetric monodromy can be witnessed by Hodge theory, and therefore be applied to similar equivariant enrichments of classical enumerative problems.
@@ 27 $[w,z,-z,-w]$
Given the labeling of the lines on the Fermat as in data A.1, the Galois group $W(E_{6})$ is given by
@@
The generators $\sigma_{1},\sigma_{2},\tau_{1},\tau_{2}\in W(E_{6})$ from 5.10 are given by the following permutations:

### v2

@@ 5. Calculating the monodromy group
In this section, we will determine the monodromy group of the cover $\widetilde{\mathcal{Y}^{\mathrm{sm}}}\to\mathcal{Y}^{\mathrm{sm}}$ by a combination of classical, moduli-theoretic, and computational techniques. We begin with the following basic fact:
@@
5.1. Certified tracking
In order to gain some insight into the structure of the symmetric monodromy group, we wish to witness the existence of certain elements by lifting explicit loops in the parameter space. In conversations with T. Brysiewicz, working with his Pandora software [Bry24], we were able to generate strong computational evidence towards the structure of the monodromy group.
While homotopy continuation software can generate strong evidence towards a computation, more refined algorithms are needed to turn these computations into proof. At each stage of tracking solutions along a one-parameter path, a guarantee is needed that paths don’t collide, and therefore that the computed permutation is indeed correct. These more sophisticated (and time-costly) methods are called certified tracking algorithms. Recent work of T. Duff and K. Lee provides algorithms which, among other things, are applicable for certifying computations in monodromy, bridging the gap between computation and proof [DL24, Theorem 1]. In conversations with Lee, their software is able to mathematically certify Theorem 1.3:
Theorem 5.5 (Numerical certification).
The monodromy group of lines on symmetric cubic surfaces is isomorphic to the Klein 4-group, and centralizes $S_{4}$ in $W(E_{6})$.
Setup.
To determine the monodromy group, it suffices to obtain generators for $\pi_{1}(\mathcal{Y}^{\mathrm{sm}})$ and then run certified tracking algorithms on their fibers. The discriminant locus of cubic surfaces, when intersected with the symmetric locus, factors into a product of four irreducible components, three of which are real lines and one of which is a cubic polynomial. Taking loops around each of these and lifting gives the desired result. Finally to identify $W(E_{6})$ as a subgroup of $S_{27}$, we must label the numerical solutions for lines at our basepoint with explicit equations for the lines on the Fermat cubic surface. ∎
Remark 5.6 (On this monodromy group).
  1. (1)
    There are further geometric constraints on monodromy that fail to completely pin down the group — for instance, the tritangent on $S_{4}$-symmetric cubic surfaces which is stabilized by the dihedral group $D_{8}$ [Bra24, Theorem 1.2] is fixed pointwise under any path in the smooth symmetric locus. This implies the monodromy group is contained in its stabilizer in $W(E_{6})$, a subgroup of order $192$ (see 5.11 below). The intersection of this subgroup with the normalizer of $S_{4}$ has order $16$, and contains the honest monodromy group properly. However, knowing the stabilizer of this tritangent will be useful when determining the connected components of the symmetric $27$ lines cover $\widetilde{\mathcal{Y}}\to\mathcal{Y}$, which we do in the next section.
  2. (2)
    The short exact sequence Equation 4 admits an explanation using the language of stacks in upcoming work of A. Landi [Lan25]. In this work, Landi resolves various equivariant monodromy problems by defining new stacks of equivariant objects. In particular, he recovers this monodromy calculation with completely different methods.
5.2. The incidence variety of 27 lines over the symmetric locus
Now that we have determined the monodromy group of the cover $\widetilde{\mathcal{Y}^{\mathrm{sm}}}\to\mathcal{Y}^{\mathrm{sm}}$ is $K_{4}$, we would like to say more about the topology of the space of symmetric cubic surfaces with a line, namely how this restricted cover splits into connected components. To do this, we will make explicit the action of $K_{4}$ on $\widetilde{\mathcal{Y}^{\mathrm{sm}}}$.
We first recall the following result of the first named author, which proves that the $S_{4}$-action on the $27$ lines is independent of the choice of smooth symmetric cubic surface:
@@ 5.2. The incidence variety of 27 lines over the symmetric locus
Once labels are fixed on the 27 lines, we can construct $W(E_{6})$ as a permutation group, given as the adjacency-preserving permutations of the 27 lines. As a subgroup of $S_{27}$ with the labeling of the lines coming from the Fermat cubic surface, the generators for $W=W(E_{6})$ are listed in data A.2.
@@
The incidence variety of 27 lines restricted to the symmetric locus $\widetilde{\mathcal{Y}^{\mathrm{sm}}}$ has 12 connected components. Explicitly as a $K_{4}$-set, the fiber over any symmetric cubic surface is of the form
@@
To conclude, we have an equivariant parameter space analog of 2.8:
@@
The space $\widetilde{\mathcal{Y}^{\mathrm{sm}}}$ is a naturally a (disconnected) complex manifold. Moreover, the equivariant monodromy representation
$\pi_{1}(\mathcal{Y}^{\mathrm{sm}},X)\to\operatorname{Aut}_{S_{4}}(H^{2}(X,\mathbb{Z}),\eta_{X})\cong S_{4}\times K_{4}$
is not surjective and has image isomorphic to the Klein $4$-group $K_{4}$.
@@ 6.3. Our groups are not reflection groups
Suppose for the sake of contradiction that $K_{4}$ was a reflection subgroup; it would then take on the form of $W(A_{1})\times W(A_{1})$. Since each of the generators $s_{i}$ act on the 27 lines as a product of six disjoint transpositions, there are two nontrivial elements of $W(A_{1})\times W(A_{1})$ that are the product of six disjoint transpositions. However, computations in GAP (using data A.3) tell us that the symmetric monodromy group only has one element that is the product of six disjoint transpositions, a contradiction. ∎
@@ 6.4. Symmetric monodromy in the projective orthogonal groups
Recall that the set of six skew lines $\{\ell_{1},\ell_{3},\ell_{10},\ell_{11},\ell_{16},\ell_{22}\}$ determine a marking of the homology of $S$, where each line corresponds to the homology classes of orthogonal $(-1)$-exceptional curves $e_{1},e_{2},e_{3},e_{4},e_{5},e_{6}$ on $S$. Using data A.3, we can calculate how the monodromy group $K_{4}$ acts on the exceptional $(-1)$-curves, which in turns explicitly determines how $K_{4}$ acts on the $E_{6}$ lattice. Then by passing to the root lattice quotient used in the proof of the exceptional isomorphism outlined in 3.3, this $K_{4}$ projects to the symmetric monodromy group $K_{4}$ inside of $\operatorname{PO}(4,1,\mathbb{F}_{3})$.
@@
As Beauville remarks [Bea09, pg. 19], what makes this difficult is that it is unknown how to produce a marking of a cubic surface from a framing of the corresponding cyclic cubic 3-fold. A resolution to this problem would shed further light on symmetric monodromy can be witnessed by Hodge theory, and therefore could be applied to similar equivariant enrichments of classical enumerative problems.
@@ 27 $[w,z,-z,-w]$
Given the labeling of the lines on the Fermat as in data A.1, the Galois group $W(E_{6})$ can be maniputaled in GAP ([GAP24]) by
@@
The generators $\sigma_{1},\sigma_{2},\tau_{1},\tau_{2}\in W(E_{6})$ from 5.12 are given by the following permutations: